# A fully relativistic treatment of EMRIs in collisionless environments

### v1

@@ Abstract
Future mHz gravitational wave (GW) interferometers will precisely probe massive black hole environments, such as accretion discs, cold dark matter overdensities, and clouds of ultralight bosons, as long as we can accurately model the dephasing they induce on the waveform of extreme mass-ratio inspirals (EMRIs). Most existing models rely on extrapolations from Newtonian results to model the interaction of the small black hole in an EMRI system with the environment surrounding the massive black hole. Here, we present a fully relativistic formalism to model such interaction with collisionless environments, focusing on the case of cold dark matter overdensities, like ‘spikes’ and ‘mounds’. We implement our new formalism in the FastEMRIWaveforms framework, and show that the resulting waveforms are significantly different from those based on a Newtonian treatment of environmental effects. Our results indicate that a fully relativistic treatment is essential to capture the environmental dephasing of GW signals from EMRIs accurately.
@@
Introduction.— Next-generation GW detectors like LISA [1], TianQin [2], and DECIGO [3] will allow us to probe the environments where compact binaries form, evolve, and merge [4, 5, 6, 7, 8]. In particular, these mHz detectors will allow us to track individual extreme mass-ratio inspirals (EMRIs) over hundreds of thousands of orbital cycles, spanning months to years, opening up new opportunities for astrophysics [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] and fundamental physics [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 33, 42, 43, 44, 45, 46, 47, 36, 48, 49, 50, 51]. The scientific potential of EMRIs hinges critically on the accuracy of their theoretical modeling. While these systems are inherently relativistic [52], their interactions with surrounding matter are often approximated using Newtonian extrapolations, sometimes supplemented by special relativity corrections (e.g., [53, 54, 55]); a few notable exceptions are, e.g., [56, 57] for accretion disks, [41, 33, 42] for ultralight bosons, and [58] for non-gravitational forces. The inaccurate modeling of environmental effects might lead to false violations of GR [59], systematic biases [22], or, in extreme cases, to missing signals [22].
In this Letter, we develop a relativistic framework to model environmental effects on EMRIs in collisionless media. We then focus on dark matter (DM) spikes as a key example, assessing the significance of relativistic effects in such systems. Our results demonstrate that a fully relativistic treatment is necessary to accurately model these interactions. We adopt geometric units ($c=G=1$) and use the mostly positive metric signature. The symbol $\overset{*}{=}$ denotes evaluation in a particular frame.
@@
Adiabatic evolution.— The adiabatic modeling of EMRIs [60, 52] relies on a two-timescale expansion [61, 62], which separates the short orbital timescale, $T_{\rm orb}\sim M$, over which dissipative effects are evaluated, from the long radiation-reaction timescale, $T_{\rm rad}\sim M/q$, over which the orbit evolves; here, $M$ is the mass of the primary BH, and $q\equiv m/M\ll 1$ is the system’s mass ratio. In this framework, the system is described by a sequence of geodesics evolving at a rate determined by the orbit-averaged dissipative fluxes of GW radiation. The adiabatic approximation is equivalent to the leading-order dissipative effects of the first-order gravitational self-force [63, 64], neglecting conservative effects and (averaged-out) higher-order dissipative effects [65].
@@
We build on this to develop a relativistic formalism to account for the effects of collisionless environments on EMRIs at adiabatic order (similarly to the approach taken in [52] for clouds of ultralight scalars). We consider a background spacetime described by the Kerr metric, neglecting the gravitational backreaction of the environment, which is tiny around EMRIs for most astrophysical systems, as both the enclosed mass within the orbit and the environment compactness are minuscule (more on that later). A geodesic $\xi$, with 4-velocity $\bm{u}\equiv\mathrm{d}\xi/\mathrm{d}\tau$ and proper time $\tau$, is uniquely determined by its energy, angular momentum around the axis of symmetry (i.e. the primary’s spin axis), and Carter’s constant defined by [^1] To completely fix the geodesic a point $\xi(\tau_{0})$ must be supplemented.
@@
The flux of GW radiation to infinity and down the primary’s event horizon can be computed by solving the Teukolsky equation, as detailed in [67, 68, 69]. Their contribution to the secular evolution of the energy and angular momentum along the primary’s spin can be found directly from global conservation laws and, while more indirectly, the evolution of the Carter’s constant can also be determined by similar techniques [70, 60, 71, 72]. A 4-acceleration $\bm{a}$ defined over $\xi$ describing the interaction of the companion with a local environment results in the additional contribution
@@
where the dot denotes the derivative with respect to the Boyer-Lindquist time $t$. These, together with the contribution of the GW fluxes, can then be orbit-averaged to compute the adiabatic evolution of the EMRI. The key question is then: how to compute the 4-acceleration from a (local) interaction with an environment within GR?
Collisionless environment.— We consider a kinetically supported collisionless medium, under the gravitational influence of a supermassive BH—this could be, for instance, a cold DM overdensity [73, 74, 75, 76] or a radiatively inefficient accretion disk [77, 78, 79]. Such media is described by a distribution function $F(x,\bm{p})$, representing the invariant number $\mathrm{d}N$ of particles contained in a spacelike volume element $\mathrm{d}V$ at $x$, and whose 4-momenta $\bm{p}$ lie within the corresponding mass-shell 3-surface element $\mathrm{d}P$ in momentum space, which is constrained by $\bm{p}\cdot\bm{p}=-\mu^{2}$, with $\mu$ the rest mass of the particles. This distribution function is independent of $\mathrm{d}V$, when defined through $\mathrm{d}N\equiv F(x,\bm{p})(-\bm{p}\cdot\bm{n})\mathrm{d}V\mathrm{d}P$, where $\bm{n}$ is the future timelike unit normal to $\mathrm{d}V$ [80]. In the following, it will be convenient to introduce (local) normal coordinates at $x$ associated to the orthonormal frame $(\bm{u},\bm{e}_{a})$. In the induced coordinates in momentum space, the volume element is $\mathrm{d}P=(1/p^{0})\,\mathrm{d}p^{1}\mathrm{d}p^{2}\mathrm{d}p^{3}$.
@@
The number of particles per phase space volume in the companion’s local plane of simultaneity (i.e., $\bm{n}=\bm{u}$) is
@@
where, for later convenience, we changed coordinates in the momentum space to $v^{a}$, defined by $p^{a}\overset{*}{=}\mu\gamma_{v}v^{a}$, with $\gamma_{v}~{}\equiv~{}(1-v^{2})^{-\nicefrac{{1}}{{2}}}$ and $v\equiv(v^{a}v_{a})^{\nicefrac{{1}}{{2}}}$. The function $f(\mathcal{E},\mathcal{C},\mathcal{L}_{z})$ depends on the type of environment and its evolution history, and can, in certain cases, be derived from first principles. At each position $x$, it is straightforward to relate $\mathcal{E}$, $\mathcal{C}$, and $\mathcal{L}_{z}$ to $v^{i}$, fixing the spatial frame vectors as $\bm{e}_{1}\propto(\bm{\partial}_{r}+u_{r}\bm{u})$, $\bm{e}_{2}\propto(\bm{\partial}_{\theta}+u_{\theta}\bm{u})$, and $\bm{e}_{3}\propto(\bm{\partial}_{\varphi}+u_{\varphi}\bm{u})$, where $(r,\theta,\varphi)$ are Boyer-Lindquist coordinates. For quasi-circular EMRIs in Schwarzschild ($r=R$, $\theta=\pi/2$, $u_{\theta}=u_{r}=0$):
$\begin{gathered}\frac{\mathcal{E}}{\varepsilon}\overset{*}{=}\gamma_{v}\frac{1+v_{3}(l_{z}/R)}{\sqrt{1+(l_{z}/R)^{2}}}\,,\qquad\mathcal{C}^{\nicefrac{{1}}{{2}}}\overset{*}{=}\gamma_{v}R\,|v_{2}|\,,\\ \frac{\mathcal{L}_{z}}{l_{z}}\overset{*}{=}\gamma_{v}\left[1+v_{3}\sqrt{1+(R/l_{z})^{2}}\right]\,,\end{gathered}$ (4)
with $\varepsilon=\frac{R-2M}{\sqrt{R(R-3M)}}$ and $|l_{z}|=\frac{R\sqrt{M}}{\sqrt{R-3M}}$, the energy and angular momentum of the companion normalized by its rest mass.
Interaction with environment.— In Fermi normal coordinates $(\tau,x^{a})$ adapted to $\xi(I)$, with $I\equiv(\tau_{1},\tau_{1}+\delta\tau)$, the tidal field from the primary BH vanishes at leading order in a neighborhood around $\xi(I)$ [82]. Even so, when we ”zoom in” close enough to the secondary object, its gravitational degrees of freedom become noticeable—this is often called the body zone [64]; we expect the linear size of this region to be roughly of the order of the Hill radius, i.e., $x\lesssim x_{\rm H}\equiv R\,(q/3)^{\nicefrac{{1}}{{3}}}$, for a circular orbit of radius $R$ and a mass-ratio $q\equiv m/M\ll 1$. The extreme mass-ratio condition is crucial to establish the separation of scales of the problem, which we will now explore.
For EMRIs immersed in incoherent collisionless media, we expect most of the energy-momentum exchange to happen locally inside the companion’s body zone. Here, we will consider the companion to be a non-spinning BH interacting with the environment through gravity, but spin effects (e.g., [83, 84]) or extra couplings to matter may be added to our framework. In the body zone, the metric is well approximated by the Schwarzschild background (here in isotropic coordinates approaching asymptotically the Fermi normal coordinates) [^2] We assume that the 4-acceleration is small enough that its effects are negligible in the body zone.
@@
The 4-acceleration arising from the gravitational interaction with the environment is (by definition)
@@
The last expression can be easily evaluated in Fermi normal coordinates, using known results for particles/fields scattering off a Schwarzschild BH. The linear momentum exchange is (in Fermi normal coordinates)
@@
with $\bm{p}_{\perp}\overset{*}{=}\mu\gamma_{v}(0,v^{a})$, $\mathrm{d}N/(\mathrm{d}V\mathrm{d}^{3}v^{i})$ as in (3), and $\sigma_{\rm acc}$ and $\sigma_{\rm scat}$, respectively, the accretion and gravitational scattering cross-sections. In the last expression, we used the isotropy of Schwarzschild. [^3] The companion’s spin (or asphericity) will increase the complexity of the problem, but does not pose an obstacle to our approach. The integral in $\mathrm{d}v^{3}\equiv\mathrm{d}v^{1}\mathrm{d}v^{2}\mathrm{d}v^{3}$ will generally have to be performed numerically.
Finally, as an example, using the same frame of the last section, the rate of change of the specific energy and angular momentum of a (equatorial) quasi-circular EMRI in Schwarzschild will be (from Eqs. (8))
$\begin{gathered}\frac{\dot{\varepsilon}_{e}}{\varepsilon}\overset{*}{=}\frac{a_{3}\left(l_{z}/R\right)}{\sqrt{1-\frac{3M}{R}}\sqrt{1+(l_{z}/R)^{2}}}\,,\quad\dot{C}_{e}^{\nicefrac{{1}}{{2}}}\overset{*}{=}\frac{R\,|a_{2}|}{\sqrt{1-\frac{3M}{R}}}\,,\\ \frac{\dot{l}_{z,e}}{|l_{z}|}\overset{*}{=}\frac{a_{3}}{\sqrt{1-\frac{3M}{R}}}\sqrt{(R/l_{z})^{2}+1}\,.\end{gathered}$ (8)
The orbit is maintained equatorial if $a^{2}$ (or its orbit-average) vanishes. In particular, this is the case if the environment’s distribution function is symmetric with respect to the equatorial plane.
The careful reader may have noticed that the frame $\bm{e}_{a}$ chosen in the previous section is not parallel transported along $\xi$. This means that, in Fermi normal coordinates, the expressions (4)—which we use to compute the integral over the distribution function in (7)—are strictly valid only at a given instant $\tau_{1}$. Nevertheless, they remain a good approximation for an interval $\delta\tau\ll R^{2}/|l_{z}|$, the timescale for the parallel transported frame to change significantly from the one we use. As the gravitational scattering happens over a very much shorter timescale $\delta\tau_{\rm int}\lesssim(x_{\rm H}/R)\sqrt{1+(l_{z}/R)^{2}}(R^{2}/|l_{z}|)\ll(R^{2}/|l_{z}|)$, [^4] We used $\delta\tau_{\rm int}\lesssim x_{\rm H}/|\braket{v^{3}}|$, assuming $\braket{\mathcal{L}_{z}}\approx 0$; but the argument holds as long as $|\braket{v^{3}}|\sim|l_{z}|/R$. we can to a good approximation use (4) to evaluate (7).
@@
Cold dark matter overdensity.— For concreteness, we focus on EMRIs in the presence of cold and collisionless DM. The density and velocity distribution of DM particles around BHs has been extensively discussed in the literature (see [6] for a recent review). If BHs grow adiabatically from an infinitesimal seed at the center of a DM halo with a power-law density profile, DM is expected to develop a dense DM ‘spike’ characterized by a density profile which is again a power-law, but with a steeper slope [85, 73]. In practice, the finite mass and size of BH seeds imply that within a radius equal to the size of the star that eventually collapses to the BH seed, the DM distribution is predicted to form an overdensity shallower than a DM ‘spike’, referred to as DM ‘mound’ [76]. A fully relativistic treatment of the evolution of the DM distribution function also leads to a substantial deviation of the DM density profile from the case of a Newtonian DM spike, both in the case of a Schwarzschild [74] and a Kerr BH [75].
@@
We consider a relativistic DM spike around a massive Schwarzschild BH, and derive the distribution $f(\mathcal{E},\mathcal{C},\mathcal{L}_{z})$ from adiabatic invariants as in [74]. [^5] As noted in footnote 3 of [75], there is a $(1-2M/r)^{-1}$ factor missing in the Schwarzschild’s radial invariant in Eq. (3.19) of [74]. We assume a Milky Way (MW) like galaxy, assuming an isotropic DM halo described by a Hernquist density profile [86] with mass $M_{\rm h}=10^{12}M_{\odot}$ and scale length $a=20\,{\rm kpc}$, and a supermassive BH with mass $M=10^{6}\,M_{\odot}$. The resulting DM spike profile has a peak rest mass density of $2.4\times 10^{19}\,\mathrm{GeV/cm^{3}}$ at $r\approx 7.7M$, as measured by stationary observers with 4-velocity $\bm{n}=(-g_{tt})^{-\nicefrac{{1}}{{2}}}\bm{\partial}_{t}$.
@@
Particle dark matter.— The process of collisionless DM scattering off a Schwarzschild BH companion is determined completely by the spacetime geodesic structure. As shown in the Appendix B of [87], the cross-sections entering in the integrand of (7) are $\sigma_{\rm acc}=\pi b_{\rm cr}^{2}(v)$, and $\sigma_{\rm scat}~{}=~{}2\pi\int_{b>b_{\rm cr}}\mathrm{d}b\,b[m\cos[\phi_{\infty}(b,v)]/b_{\rm gr}(v)]^{2}$, where [87]
[…]
$b_{\rm gr}=\frac{m}{v^{2}}\left(1+v^{2}\right)\,,\qquad\pi-2\phi_{\infty}=\;\text{deflection angle}\,.$
The asymptotic true anomaly $\phi_{\infty}$ can be expressed in terms of elliptic integrals of the first kind [88, 87].
@@
Figure 1: Energy loss per orbital cycle as a function of binary separation. Solid lines correspond to the fully relativistic model, the dashed to the Newtonian, and the dotted to including heuristic relativistic corrections as in [54], for an EMRI with mass-ratio $q=10^{-5}$ in a DM spike of a MW-like galaxy. An EMRI at $R=12M$ will merge after $\sim 4$ years.
@@
The previous expression can be made more explicit by replacing $\sigma_{\rm scat}\equiv 4\pi\Lambda\,b_{\rm gr}^{2}$, introducing the Coulomb logarithm (using an infrared cutoff $b_{\rm max}\sim x_{\rm H}$)
@@
where $b_{\rm min}$ is chosen freely, provided $b_{\rm cr}\ll b_{\rm min}\leq b_{\rm max}$. The function $\chi$ is monotonically increasing and $O(1)$ for $v\gtrsim 10^{-1}$. It is accurately approximated by a simple analytic expression given in the Supplemental Material.
@@
In both models, we keep the relativistic expressions (4) in the particle distribution function, to preserve the loss cone of the DM distribution; a Newtonian modeling of the loss cone would result in a depleted region $r<8M$ [73] (as opposed to the actual $r<4M$ [74]). We also keep $b_{\rm cr}(v)$ to compute the effect of accretion, despite that gravitational capture is a relativistic phenomenon. We neglect the relativistic factors $\gamma_{v}$ in the incoming momentum $\bm{p}_{\perp}$, $(1+v^{2})$ in the scattering impact parameter $b_{\rm gr}$, $\chi$ in the Coulomb logarithm $\Lambda$, the strong-field terms in the specific torque in (8), and we use $\mathrm{d}N/(\mathrm{d}V\mathrm{d}^{3}v^{i})$ as measured by the stationary observers [$\bm{n}=(-g_{tt})^{-\nicefrac{{1}}{{2}}}\bm{\partial}_{t}$] in the integration (7). The latter (multiplied by the rest mass $\mu$) coincides with the mass density obtained in [74], which has been used in previous models of environmental effects.
For comparison, we also consider a phenomenological model of the relativistic corrections motivated by [54] (labeled ‘Phenom.’). This includes a multiplicative factor $\gamma_{u}^{2}\equiv 1/(1-u^{2})$, with $u\equiv\Omega R$, to the dynamical friction and accretion torques, as an attempt to account for the volume contraction and the relativistic correction to the particles momentum from the companion’s perspective, and also a multiplicative factor $(1+u^{2})^{2}$ to the dynamical friction torque, from relativistic corrections to $b_{\rm gr}^{2}$ (cf. previous paragraph). The mismatch with the fully relativistic model (shown in Fig. 1) can be understood from the non-inclusion of strong-field effects from both the supermassive and companion BHs, but also from taking the (heuristic) relativistic effects out of the integral over the velocity space (7) (see Supplemental Material).
@@
Impact of relativistic effects.— We quantify the difference between two waveform-modeling frameworks using the ‘faithfulness’ integral $\mathcal{F}\equiv\left<a|b\right>/\left(\rho_{a}\rho_{b}\right)$ [89] and its complement, the ‘mismatch’ $\mathcal{M}=1-\mathcal{F}$, where the noise-weighted inner product between two real waveforms $a(t)$ and $b(t)$ is given in the frequency-domain by
$a|b\right>\equiv 4\,\mathrm{Re}\int_{f_{\mathrm{min}}}^{\infty}\frac{a^{*}\left(f\right)b\!\left(f\right)}{S_{n}\left(f\right)}\,\mathrm{d}f\,,$ (9)
@@
with $S_{n}\left(f\right)$ the one-sided LISA power-spectral density (from within FEW), and $\rho_{a}~{}\!\!\!\equiv~{}\!\!\!\sqrt{\left<a|a\right>}$ the optimal signal-to-noise-ratio (SNR). We evaluate Eq. (9) numerically from $f_{\mathrm{min}}~{}\!\!=~{}\!\!10^{-5}$ Hz up to the Nyquist frequency, and maximize the result over phase and time offsets as in [30]. We obtain the EMRI waveforms in each case by incorporating Eqs. 8 into the FastEMRIWaveforms (FEW) framework [90, 91], initializing binaries 4 years before merger.

### v2

# Fully Relativistic Treatment of Extreme Mass-Ratio Inspirals in Collisionless Environments

@@ Abstract
Future mHz gravitational wave (GW) interferometers will precisely probe massive black hole environments, such as accretion disks, cold dark matter overdensities, and clouds of ultralight bosons, as long as we can accurately model the dephasing they induce on the waveform of extreme mass-ratio inspirals (EMRIs). Most existing models rely on extrapolations from Newtonian results to model the interaction of the small black hole in an EMRI system with the environment surrounding the massive black hole. Here, we present a fully relativistic formalism to model such interaction with collisionless environments, focusing on the case of cold dark matter overdensities, like “spikes” and “mounds”. We implement our new formalism in the FastEMRIWaveforms framework, and show that the resulting waveforms are significantly different from those based on a Newtonian treatment of environmental effects. Our results indicate that a fully relativistic treatment is essential to capture the environmental dephasing of GW signals from EMRIs accurately.
@@
Introduction.— Next-generation GW detectors like LISA [1], TianQin [2], and DECIGO [3] will allow us to probe the environments where compact binaries form, evolve, and merge [4, 5, 6, 7, 8]. In particular, these mHz detectors will allow us to track individual extreme mass-ratio inspirals (EMRIs) over hundreds of thousands of orbital cycles, spanning months to years, opening up new opportunities for astrophysics [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] and fundamental physics [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 35, 44, 45, 46, 47, 48, 49, 38, 50, 51, 52, 53]. The scientific potential of EMRIs hinges critically on the accuracy of their theoretical modeling. While these systems are inherently relativistic [54], their interactions with surrounding matter are often approximated using Newtonian extrapolations, sometimes supplemented by special relativity corrections (e.g., [55, 56, 57]); a few notable exceptions are, e.g., [58, 59, 25] for accretion disks, [43, 35, 44] for ultralight bosons, and [60] for non-gravitational forces. The inaccurate modeling of environmental effects might lead to false violations of general relativity [61], systematic biases [22], or, in extreme cases, to missing signals [22].
[…]
@@
Adiabatic evolution.— The adiabatic modeling of EMRIs [62, 54] relies on a two-timescale expansion [63, 64], which separates the short orbital timescale, $T_{\rm orb}\sim M$, over which dissipative effects are evaluated, from the long radiation-reaction timescale, $T_{\rm rad}\sim M/q$, over which the orbit evolves; here, $M$ is the mass of the primary black hole (BH), and $q\equiv m/M\ll 1$ is the system’s mass ratio. In this framework, the system is described by a sequence of geodesics evolving at a rate determined by the orbit-averaged dissipative fluxes of GW radiation. The adiabatic approximation is equivalent to the leading-order dissipative effects of the first-order gravitational self-force [65, 66], neglecting conservative effects and (averaged-out) higher-order dissipative effects [67].
@@
We build on this to develop a relativistic formalism to account for the effects of collisionless environments on EMRIs at adiabatic order (similarly to the approach taken in [54] for clouds of ultralight scalars). We consider a background spacetime described by the Kerr metric, neglecting the gravitational backreaction of the environment, which is tiny around EMRIs for most astrophysical systems, as both the enclosed mass within the orbit and the environment’s compactness are minuscule (more on that later). A geodesic $\xi$, with four-velocity $\bm{u}\equiv\mathrm{d}\xi/\mathrm{d}\tau$ and proper time $\tau$, is uniquely determined by its energy, angular momentum around the axis of symmetry (i.e. the primary’s spin axis), and Carter’s constant defined by [^1] To completely fix the geodesic a point $\xi(\tau_{0})$ must be supplemented.
@@
The flux of GW radiation to infinity and down the primary’s event horizon can be computed by solving the Teukolsky equation, as detailed in [69, 70, 71]. Their contribution to the secular evolution of the energy and angular momentum along the primary’s spin can be found directly from global conservation laws and, while more indirectly, the evolution of the Carter’s constant can also be determined by similar techniques [72, 62, 73, 74]. A four-acceleration $\bm{a}$ defined over $\xi$ describing the interaction of the companion with a local environment results in the additional contribution
@@
where the dot denotes the derivative with respect to the Boyer-Lindquist time $t$. These, together with the contribution of the GW fluxes, can then be orbit-averaged to compute the adiabatic evolution of the EMRI. The key question is then: how to compute the four-acceleration from a (local) interaction with an environment within general relativity?
Collisionless environment.— We consider a kinetically supported collisionless medium, under the gravitational influence of a supermassive BH—this could be, for instance, a cold DM overdensity [75, 76, 77, 78] or a radiatively inefficient accretion disk [79, 80, 81]. Such media is described by a distribution function $F(x,\bm{p})$, representing the invariant number $\mathrm{d}N$ of particles contained in a spacelike volume element $\mathrm{d}V$ at $x$, and whose four-momenta $\bm{p}$ lie within the corresponding mass-shell three-surface element $\mathrm{d}P$ in momentum space, which is constrained by $\bm{p}\cdot\bm{p}=-\mu^{2}$, with $\mu$ the rest mass of the particles. This distribution function is independent of $\mathrm{d}V$, when defined through $\mathrm{d}N\equiv F(x,\bm{p})(-\bm{p}\cdot\bm{n})\mathrm{d}V\mathrm{d}P$, where $\bm{n}$ is the future timelike unit normal to $\mathrm{d}V$ [82]. In the following, it will be convenient to introduce (local) normal coordinates at $x$ associated to the orthonormal frame $(\bm{u},\bm{e}_{a})$. In the induced coordinates in momentum space, the volume element is $\mathrm{d}P=(1/p^{0})\,\mathrm{d}p^{1}\mathrm{d}p^{2}\mathrm{d}p^{3}$.
@@
The number of particles per phase-space volume in the companion’s local plane of simultaneity (i.e., $\bm{n}=\bm{u}$) is
@@
where, for later convenience, we changed coordinates in the momentum space to $v^{a}$, defined by $p^{a}\overset{*}{=}\mu\gamma_{v}v^{a}$, with $\gamma_{v}~\equiv~(1-v^{2})^{-\nicefrac{{1}}{{2}}}$ and $v\equiv(v^{a}v_{a})^{\nicefrac{{1}}{{2}}}$. The function $f(\mathcal{E},\mathcal{C},\mathcal{L}_{z})$ depends on the type of environment and its evolution history, and can, in certain cases, be derived from first principles. At each position $x$, it is straightforward to relate $\mathcal{E}$, $\mathcal{C}$, and $\mathcal{L}_{z}$ to $v^{i}$, fixing the spatial frame vectors as $\bm{e}_{1}\propto(\bm{\partial}_{r}+u_{r}\bm{u})$, $\bm{e}_{2}\propto(\bm{\partial}_{\theta}+u_{\theta}\bm{u})$, and $\bm{e}_{3}\propto(\bm{\partial}_{\varphi}+u_{\varphi}\bm{u})$, where $(r,\theta,\varphi)$ are Boyer-Lindquist coordinates. The explicit expressions for quasi-circular EMRIs in Schwarzschild are given in the End Matter.
Interaction with environment.— In Fermi normal coordinates $(\tau,x^{a})$ adapted to $\xi(I)$, with $I\equiv(\tau_{1},\tau_{1}+\delta\tau)$, the tidal field from the primary BH vanishes at leading order in a neighborhood around $\xi(I)$ [84]. Even so, when we ”zoom in” close enough to the secondary object, its gravitational degrees of freedom become noticeable—this is often called the body zone [66]; we expect the linear size of this region to be roughly of the order of the Hill radius, i.e., $x\lesssim x_{\rm H}\equiv R\,(q/3)^{\nicefrac{{1}}{{3}}}$, for a circular orbit of radius $R$ and a mass-ratio $q\equiv m/M\ll 1$. The extreme mass-ratio condition is crucial to establishing the separation of scales of the problem, which we will now explore.
For EMRIs immersed in incoherent collisionless media, we expect most of the energy-momentum exchange to happen locally inside the companion’s body zone. Here, we will consider the companion to be a non-spinning BH interacting with the environment through gravity, but spin effects (e.g., [85, 86]) or extra couplings to matter may be added to our framework. In the body zone, the metric is well approximated by the Schwarzschild background (here in isotropic coordinates approaching asymptotically the Fermi normal coordinates) [^2] We assume that the four-acceleration is small enough that its effects are negligible in the body zone.
@@
The four-acceleration arising from the gravitational interaction with the environment is (by definition)
@@
The last expression can be easily evaluated in Fermi normal coordinates, using known results for particles scattering off a Schwarzschild BH. The linear momentum exchange is (in Fermi normal coordinates)
@@
with $\bm{p}_{\perp}\overset{*}{=}\mu\gamma_{v}(0,v^{a})$, $\mathrm{d}N/(\mathrm{d}V\mathrm{d}^{3}v^{i})$ as in (3), and $\sigma_{\rm acc}$ and $\sigma_{\rm scat}$, respectively, the accretion (acc) and gravitational scattering (scat) cross sections. In the last expression, we used the isotropy of Schwarzschild. [^3] The companion’s spin (or asphericity) will increase the complexity of the problem, but does not pose an obstacle to our approach. The integral in $\mathrm{d}v^{3}\equiv\mathrm{d}v^{1}\mathrm{d}v^{2}\mathrm{d}v^{3}$ will generally have to be performed numerically.
The acceleration computed in Fermi normal coordinates adapted to the companion has then to be related to the rate of change of the conserved quantities describing its motion. Explicit expressions for quasi-circular EMRIs in Schwarzschild are given in the End Matter.
@@
Cold dark matter overdensity.— For concreteness, we focus on EMRIs in the presence of cold and collisionless DM. The density and velocity distribution of DM particles around BHs has been extensively discussed in the literature (see [6] for a recent review). If BHs grow adiabatically from an infinitesimal seed at the center of a DM halo with a power law density profile, DM is expected to develop a dense DM “spike” characterized by a density profile which is again a power law, but with a steeper slope [87, 75]. In practice, the finite mass and size of BH seeds imply that within a radius equal to the size of the star that eventually collapses to the BH seed, the DM distribution is predicted to form an overdensity shallower than a DM spike, referred to as DM “mound” [78]. A fully relativistic treatment of the evolution of the DM distribution function also leads to a substantial deviation of the DM density profile from the case of a Newtonian DM spike, both in the case of a Schwarzschild [76] and a Kerr BH [77].
@@
We consider a relativistic DM spike around a massive Schwarzschild BH, and derive the distribution $f(\mathcal{E},\mathcal{C},\mathcal{L}_{z})$ from adiabatic invariants as in [76]. [^4] As noted in footnote 3 of [77], there is a $(1-2M/r)^{-1}$ factor missing in the Schwarzschild’s radial invariant in Eq. (3.19) of [76]. We consider a Milky-Way-like galaxy, assuming an isotropic host DM halo described by a Hernquist density profile [88] with mass $M_{\rm h}=10^{12}M_{\odot}$ and scale length $a=20\,{\rm kpc}$, and a supermassive BH with mass $M=10^{6}\,M_{\odot}$. The resulting DM spike profile has a peak rest mass density of $2.4\times 10^{19}\,\mathrm{GeV/cm^{3}}$ at $r\approx 7.7M$, as measured by stationary observers with four-velocity $\bm{n}=(-g_{tt})^{-\nicefrac{{1}}{{2}}}\bm{\partial}_{t}$; this scales approximately with $\propto M^{-1.7}$ (e.g., [56]).
@@
Particle dark matter.— The process of collisionless DM scattering off a Schwarzschild BH companion is determined completely by the spacetime geodesic structure. As shown in App. B of [89], the cross sections in (6) are $\sigma_{\rm acc}=\pi b_{\rm cr}^{2}(v)$ and $\sigma_{\rm scat}=2\pi\int_{b>b_{\rm cr}}\mathrm{d}b\,b[m\cos[\phi_{\infty}(b,v)]/b_{\rm gr}(v)]^{2}$, where $b_{\rm gr}\equiv(m/v^{2})(1+v^{2})$ and the critical (capture) impact parameter, $b_{\rm cr}$, is expressed analytically in [89]. The asymptotic true anomaly $\phi_{\infty}$ can be expressed in terms of elliptic integrals of the first kind [90, 89].
@@
Figure 1: Energy loss per orbital cycle as a function of binary separation. Solid lines correspond to the fully relativistic model, the dashed to the Newtonian, and the dotted to including heuristic relativistic corrections as in [56], for an EMRI with mass-ratio $q=10^{-5}$ in a DM spike of a Milky-Way-like galaxy. An EMRI at $R=12M$ will merge after $\sim 4$ years.
@@
The scattering cross section can be made more explicit by expressing $\sigma_{\rm scat}\equiv 4\pi\Lambda\,b_{\rm gr}^{2}$, introducing the Coulomb logarithm (using an infrared cutoff $b_{\rm max}\sim x_{\rm H}$)
@@
where $b_{\rm min}$ is chosen freely, provided $b_{\rm cr}\ll b_{\rm min}\leq b_{\rm max}$. The function $\chi$ is monotonically increasing and $O(1)$ for $v\gtrsim 10^{-1}$. It is accurately approximated by a simple analytic expression given in the Supplemental Material [91].
@@
In both models, we keep the relativistic expressions (7) in the distribution function, to preserve the loss cone of the DM distribution; a Newtonian modeling of the loss cone would result in a depleted region $r<8M$ [75] (as opposed to the actual $r<4M$ [76]). We also keep $b_{\rm cr}(v)$ to compute the effect of accretion, despite that gravitational capture is a relativistic effect. We neglect the relativistic factors $\gamma_{v}$ in the incoming momentum $\bm{p}_{\perp}$, $(1+v^{2})$ in the scattering impact parameter $b_{\rm gr}$, $\chi$ in the Coulomb logarithm $\Lambda$, the strong-field terms in the specific torque in (8), and use $\mathrm{d}N/(\mathrm{d}V\mathrm{d}^{3}v^{i})$ as measured by the stationary observers $\bm{n}=(-g_{tt})^{-\nicefrac{{1}}{{2}}}\bm{\partial}_{t}$] in the integration (6). The latter (multiplied by the rest mass $\mu$) coincides with the mass density obtained in [76], which has been used in previous models of environmental effects.
For comparison, we also consider a phenomenological model of the relativistic corrections motivated by [56] (labeled “Phenom.”). This includes a multiplicative factor $\gamma_{u}^{2}\equiv 1/(1-u^{2})$, with $u\equiv\Omega R$, to the dynamical friction and accretion torques, as an attempt to account for the volume contraction and the relativistic correction to the particles momentum from the companion’s perspective, and also a multiplicative factor $(1+u^{2})^{2}$ to the dynamical friction torque, from relativistic corrections to $b_{\rm gr}^{2}$ (cf. previous paragraph). The deviation from the fully relativistic model (shown in Fig. 2) arises from neglecting strong-field effects of both the supermassive and companion BHs, and from considering the (heuristic) relativistic corrections out of the integral over the velocity space (6) (see Supplemental Material [91]).
@@
Impact of relativistic effects.— We quantify the difference between two waveform-modeling frameworks using the “faithfulness” integral $\mathcal{F}\equiv\left<a|b\right>/\left(\rho_{a}\rho_{b}\right)$ [92] and its complement, the “mismatch” $\mathcal{M}=1$-$\mathcal{F}$, where the noise-weighted inner product between two real waveforms $a(t)$ and $b(t)$ is given in frequency domain by the usual form (e.g., [93, Eq. (2.3)]), with the one-sided LISA power-spectral density taken to be the numerical Michelson-like PSD from [94], and $\rho_{a}~\!\!\!\equiv~\!\!\!\sqrt{\left<a|a\right>}$ is the optimal signal-to-noise-ratio (SNR). We evaluate the noise-weighted inner product from $f_{\mathrm{min}}~\!\!=~\!\!10^{-5}$ Hz up to the Nyquist frequency $f_{\mathrm{max}}=0.05$ Hz, and maximize the result over phase and time offsets as in [32]. We obtain the EMRI waveforms in each case by incorporating Eqs. 8 into the FastEMRIWaveforms_v1.5.4 (FEW) framework [95, 96, 97, 98], initializing binaries 4 yr before merger.